OpenStudy (pulsified333):

A coin is flipped. If a heads is flipped, then the coin is flipped 5 more times and the number of heads flipped is noted; otherwise (i.e., a tails is flipped on the initial flip), then the coin is flipped 4 more times and the result of each flip (i.e., heads or tails) is noted successively. How many possible outcomes are in the sample space of this experiment?

OpenStudy (pulsified333):

@me10000

OpenStudy (anonymous):

I don't know this sorry...

OpenStudy (pulsified333):

its ok

OpenStudy (mathmate):

Number of outcomes are there when we toss 1 time = 2^1 = 2 toss 2 times = 2^2 = 4 toss 3 times = 2^3 = 8 toss 4 times = 2^4 = 16 toss 5 times = 2^5 = 32 Then complete the following tree diagram: